# Not All Correct Answers Are Equal: Why Your Distillation Source Matters

### v1

@@ Abstract
Distillation has emerged as a practical and effective approach to enhance the reasoning capabilities of open-source language models. In this work, we conduct a large-scale empirical study on reasoning data distillation by collecting verified outputs from three state-of-the-art teacher models—AM-Thinking-v1, Qwen3-235B-A22B, and DeepSeek-R1—on a shared corpus of 1.89 million queries. We construct three parallel datasets and analyze their distributions, revealing that AM-Thinking-v1-distilled data exhibits greater token length diversity and lower perplexity. Student models trained on each dataset are evaluated on reasoning benchmarks including AIME2024, AIME2025, MATH500, and LiveCodeBench. The AM-based model consistently achieves the best performance (e.g., 84.3 on AIME2024, 72.2 on AIME2025, 98.4 on MATH500, and 65.9 on LiveCodeBench) and demonstrates adaptive output behavior—producing longer responses for harder tasks and shorter ones for simpler tasks. These findings highlight the value of high-quality, verified reasoning traces. We release the AM-Thinking-v1 and Qwen3-235B-A22B distilled datasets to support future research on open and high-performing reasoning-oriented language models. The datasets are publicly available on Hugging Face111Datasets are available on Hugging Face: AM-Thinking-v1-Distilled, AM-Qwen3-Distilled..
@@ 1 Introduction
These results highlight the practical value of large-scale, high-quality reasoning data distillation, for improving open-source LLMs. To promote further progress in the field, we release both the AM-Thinking-v1 and Qwen3-235B-A22B distilled datasets222Note that since the distillation data of DeepSeek-R1 is easily accessible, we only release the distillation data of AM-Thinking-v1 and Qwen3-235B-A22B.. We hope our work provides valuable resources and insights for the open-source community, enabling more effective reasoning-focused model development and contributing to the broader progress of reasoning research.
@@ 2.2 Data Distilling
Mathematical Reasoning: Responses were verified using a two-stage process—first with Math-Verify333https://github.com/huggingface/Math-Verify, and, if necessary, subsequently with Qwen2.5-7B-Instruct [27, 28]. Each result was assigned a binary verification score.

### v2

@@ Abstract
Distillation has emerged as a practical and effective approach to enhance the reasoning capabilities of open-source language models. In this work, we conduct a large-scale empirical study on reasoning data distillation by collecting verified outputs from three state-of-the-art teacher models—AM-Thinking-v1, Qwen3-235B-A22B, and DeepSeek-R1—on a shared corpus of 1.89 million queries. We construct three parallel datasets and analyze their distributions, revealing that AM-Thinking-v1-distilled data exhibits greater token length diversity and lower perplexity. Student models trained on each dataset are evaluated on reasoning benchmarks including AIME2024, AIME2025, MATH500, and LiveCodeBench. The model distilled from AM-Thinking-v1 consistently achieves the best performance (e.g., 84.3 on AIME2024, 72.2 on AIME2025, 98.4 on MATH500, and 65.9 on LiveCodeBench) and demonstrates adaptive output behavior—producing longer responses for harder tasks and shorter ones for simpler tasks. These findings highlight the value of high-quality, verified reasoning traces. We release the AM-Thinking-v1 and Qwen3-235B-A22B distilled datasets to support future research on open and high-performing reasoning-oriented language models. The datasets are publicly available on Hugging Face111https://huggingface.co/datasets/a-m-team/AM-Thinking-v1-Distilled..
@@ 1 Introduction
These results highlight the practical value of large-scale, high-quality reasoning data distillation, for improving open-source LLMs. To promote further progress in the field, we release both the AM-Thinking-v1 and Qwen3-235B-A22B distilled datasets444Note that since the distillation data of DeepSeek-R1 is easily accessible, we only release the distillation data of AM-Thinking-v1 and Qwen3-235B-A22B.. We hope our work provides valuable resources and insights for the open-source community, enabling more effective reasoning-focused model development and contributing to the broader progress of reasoning research.
@@ 2.2 Data Distilling
Mathematical Reasoning: Responses were verified using a two-stage process—first with Math-Verify555https://github.com/huggingface/Math-Verify, and, if necessary, subsequently with Qwen2.5-7B-Instruct [27, 28]. Each result was assigned a binary verification score.